# Transfer ABCD Matrix for Time-Varying Media and Time Crystals

### v1

@@ Abstract
This paper introduces a formal definition of the transfer ABCD parameters in time-varying electromagnetic systems. The formal definition comes after the rearrangement of the fields $D$ and $B$ at the inputs and outputs of the temporal system based on the time-varying boundary conditions. Then, we derive the ABCD parameters of a temporal transmission line, i.e., a temporal slab, and compute the associated scattering parameters (reflection and transmission coefficients). The results presented here open up an alternative way, based on network theory, to analyze multilayer temporal configurations. Moreover, we show that the ABCD parameters can be used to compute the dispersion diagram ($\omega$ vs $k$) of time crystals.
Index Terms:
Network analysis, time-varying systems, ABCD parameters, scattering parameters, S-matrix, temporal slabs, time crystal.
@@ I Introduction
Maxwell’s equations describe with precision any electric and magnetic phenomena, or the interaction between both, involving classical field theory [1]. This includes the understanding and control of electromagnetic (EM) wave propagation, radiation and wave guiding and light-matter interactions at the human and nano scales, the production of basic electronic components such as resistors, inductors, capacitors or transformers, to name a few, the use of metamaterials, and many other interesting applications [2, 3, 4, 5, 6]. Certainly, Maxwell’s equations have transcended and transformed our society.
@@
In network analysis, a microwave or photonic network can be defined as a $N$-port system, whose inputs and outputs are interconnected via a governing matrix [10]. Impedance $[Z]$ (admittance $[Y]$) matrices are widely used solutions to relate the input voltages (currents) and output currents (voltages) in the network. Similarly, the scattering matrix $[S]$ does so with incident and reflected power waves. However, in many practical scenarios, the number of ports is reduced to $N=2$ (one input and one output). In this context, the use of ABCD parameters becomes particularly relevant.
@@
Despite the well-known foundations of ABCD parameters and their vast application in time-invariant engineering scenarios, the analogous version of these parameters applied to the resolution of time-varying EM problems has not been formally established yet. In this paper, we propose a formal definition for the temporal ABCD parameters. We exploit the temporal boundary conditions of the fields $\mathbf{D}$ and $\mathbf{B}$ to rearrange the input and output of the two-port time-varying system in a way that the resulting transfer matrix can be cascaded in a similar manner to conventional ABCD parameters. Then, we extract the ABCD parameters of a temporal transmission line (a temporal slab), and connect them with the associated scattering parameters and the computation of the dispersion diagram ($\omega$ vs $k$) in time crystals.
@@ II Temporal Transfer ABCD Matrix
Fig. 1: Use of transfer ABCD parameters in multilayer spatial and temporal interfaces. In both scenarios, ABCD matrices can be used to connect the EM fields at the input and output of the system. By exploiting the continuity of tangential {$E_{t}$,$H_{t}$} (total {$D$, $B$}) fields in the spatial (temporal) problems, the fields can be rearranged such that the ABCD matrix of each layer can be multiplied to obtain the global response of the time-invariant (time-varying) system. The multilayer composite is formed by $N+2$ layers: the input layer [$(i)$, $n=0$], $N$ intermediate layers, and the output layer [$(o)$, $n=N+1$].
@@
As mentioned above, the ABCD-matrix formalism has proven to be quite efficient in characterizing cascade connections. As it is well known [10], the propagation of waves through a spatial dielectric layer of thickness $d_{n}$ admits to be represented by the following matrix
@@
where $D_{n-1},B_{n-1},D_{n},B_{n}$ are the electric/magnetic flux vectors in the $n$th-order interface. As we will see below, the result of the matrix allows us to indentify a temporal transmission line.
@@ II-A Temporal Transmission Line (Temporal Slab)
Let us consider the scenario depicted in the right panel of Fig. 1. In the scenario, a certain medium or material changes its electrical (or magnetic) properties over time at $t=t_{n-1}$. After a short period, the medium changes again its properties at $t=t_{n}$. The temporal region between $t_{n-1}$ and $t_{n}$ defines a temporal slab, a homogeneous medium that covers the whole space and is present for a limited time. Analogously to a spatial slab, a temporal slab is defined by its temporal length $T=t_{n}-t_{n-1}$, as well as its permittivity $\varepsilon_{n}$, permeability $\mu_{n}$ and the associated wave impedance $Z_{n}$. Please note that the wave impedance is defined as $Z_{n}=B_{n}/D_{n}$ in the $n$-th temporal slab.
A temporal slab can be formally described as a temporal transmission line, in analogy to the conventional spatial transmission line [10]. A temporal transmission line is an object that connects two time instants $t_{1}$ and $t_{2}$. For simplicity, we consider $t_{1}=0$ and $t_{2}=T$. Without loss of generality, the $D$ and $B$ fields can be represented at the input of the temporal transmission line in terms of forward ($+$) and backward (--) waves,
@@
Since we work with linear, isotropic and non-dispersive media, $D(t)=\varepsilon(t)E(t)$ and $B(t)=\mu(t)H(t)$ must hold. Thus, the coefficients $D^{+}$, $D^{-}$, $B^{+}$ and $B^{-}$ are related as follows
@@
After replacing eq. (6) into the expression for $D_{1}$ in (3), we realize that fields $D_{1}$, $D_{2}$ and $B_{1}$, $B_{2}$ are related as
@@
Similarly, the replacement of eq. (6) into the expression for $B_{n}$ leads to
@@
The connection between the fields at the input and at output of the temporal transmission line can be better rearranged into matrix form:
@@
thus revealing that $A^{\prime}=\cos(\omega T)$, $B^{\prime}=-\mathrm{j}\frac{1}{Z}\sin(\omega T)$, $C^{\prime}=-\mathrm{j}Z\sin(\omega T)$, and $D^{\prime}=\cos(\omega T)$ are the transfer ABCD parameters of the temporal transmission line (temporal slab). In addition, it can be rapidly checked that the network is reciprocal ($A^{\prime}D^{\prime}-B^{\prime}C^{\prime}=1$) and symmetrical ($A^{\prime}=D^{\prime}$) in terms of electric and magnetic flux densities $D,B$. This temporal or momentum reciprocity does not necessarily conduct to the reciprocity for $E,H$ (spatial reciprocity), and vice versa.
@@
In the ABCD format, the number of matrices to multiply coincides with the number of temporal slabs $N$. Notice that this entails significant differences with other matrix formalisms reported in the literature, such as that in [18]. There, the authors present a sound matrix approach that connects forward and backward $E$ waves, in a similar fashion to using scattering parameters, instead of directly working with $D$ and $B$. Since $E$ is not continuous across time interfaces, additional matrices (called matching matrices) must be added for the entire calculation. For a temporal multilayer made of $N$ slabs, the method in [18] needs to employ $12N+$ 1 matrices, in contrast to the method here presented.
@@ II-B Associated Scattering Parameters
The analogy between spatial and temporal problems gives rise to the possibility of extracting the S-parameters associated with the temporal ABCD matrix. Now, the formal definition of the scattering matrix (S-matrix) takes the $D^{\pm}$ forward/backward waves in the input/output media (1), (2). This contrasts to the spatial case, based on electric-field counterparts $E^{\pm}$. In both cases, however, there is a direct correspondence between them thanks to the constitutive relation $D^{\pm}=\varepsilon E^{\pm}$. The temporal S-matrix can be defined as:
$\begin{pmatrix}D_{2}^{-}\\ D_{2}^{+}\end{pmatrix}=\begin{pmatrix}S_{11}&S_{12}\\ S_{21}&S_{22}\end{pmatrix}\begin{pmatrix}D_{1}^{+}\\ D_{1}^{-}\end{pmatrix}\,,$ (10)
@@
$S_{11}$ $=\triangle[-A^{\prime}-B^{\prime}Z_{0}+C^{\prime}/Z_{0}+D^{\prime}]$ (11)
$S_{12}$ $=\triangle[-A^{\prime}+B^{\prime}Z_{0}-C^{\prime}/Z_{0}+D^{\prime}]$ (12)
$S_{21}$ $=\triangle[A^{\prime}-B^{\prime}Z_{0}-C^{\prime}/Z_{0}+D^{\prime}]$ (13)
$S_{22}$ $=\triangle[A^{\prime}+B^{\prime}Z_{0}+C^{\prime}/Z_{0}+D^{\prime}]$ (14)
with $\triangle=[2A^{\prime}D^{\prime}-2C^{\prime}B^{\prime}]^{-1}$ and $Z_{0}=Z^{(i)}=Z^{(o)}$. By assuming identical input/output media, the corresponding S-matrix for the electric field coincides with those in (11)-(14),
$S_{ij}=\frac{D_{2}^{\pm}}{D_{1}^{\pm}}=\frac{E_{2}^{\pm}}{E_{1}^{\pm}}\,.$ (15)
@@ II-C Considerations about Conservation of Momentum Density
[…]
@@
[…]
@@ II-D Time Crystal: Infinite Temporal Transmission Lines
with $a=0.5(\pm\sqrt{A^{\prime 2}-2A^{\prime}D^{\prime}+4B^{\prime}C^{\prime}+D^{\prime 2}}+A^{\prime}+D^{\prime})$. Eq. (22) offers a pathway to compute the dispersion diagram, $\omega$ vs $k$, in time crystals. In the above, note that $A^{\prime}D^{\prime}-B^{\prime}C^{\prime}=1$.
The advantage of the approach of the present paper is the possibility of extracting analytical expressions to compute $\omega-k$ diagrams. Other methodologies in the literature, such as that reported in [21], address the problem from a numerical perspective by using complex root-finding algorithms. Our methodology provides formal analytical expressions for $\omega(k)$, including complex-valued frequency solutions. This is a significant advantage since evanescent- and, especially, growing-field solutions manifested via the imaginary part of the frequency can not be lightly disregarded in time-varying systems [22].
@@ III-A Temporal multilayer slab
Fig. 2: Scattering parameters of an equally travel-distance multi-layered temporal structure with identical input and output media. $N=4$, $Z_{1}=Z_{0}/9$, $Z_{2}=Z_{0}/3$, and $Z^{(i)}=Z^{(o)}=Z_{0}$. Frequencies are normalized to the central frequency $f_{0}$.
@@ III-B Time Crystal
Fig. 3: Dispersion diagram, computed with the ABCD formalism (white dotted lines) and the exact numerical solution (colored plot), for a time crystal given by the temporal modulation $\varepsilon_{r}(t)=\varepsilon_{r0}\left[1+m\cos(\Omega t)\right]$. (a) Discretization of the continuous temporal modulation (blue line) into $N$ steps (black lines). Dispersion diagram in the cases: (b) $N=5$, (c) $N=30$. Parameters: $\varepsilon_{r0}=1$, $T=0.1$ ns, $m=0.7$, and $\mu_{r}=1$.
@@
[…]
@@
Finally, it is important to remark that the present method allows to compute both the real and imaginary parts of the frequency. In time crystals, imaginary frequency components refer to time decay/grow of the corresponding fields. These solutions are typical from the bandgap regions, manifesting from $k/K\in[0.4,0.55]$ in Fig. 3(b)-(c). Moreover, it is expected that the present ABCD approach can be extended to the analysis of uniform-velocity space-time crystals, in a similar fashion to was carried out in [25].

### v2

@@ Abstract
This paper introduces a formal description of the transfer ABCD parameters in time-varying electromagnetic systems. The formal description comes after the rearrangement of the electric displacement field $D$ and magnetic flux density field $B$ at the inputs and outputs of the temporal system based on the time-varying boundary conditions. Then, we derive the ABCD parameters of a temporal transmission line, i.e., a temporal slab, and compute the associated scattering parameters (reflection and transmission coefficients). The results presented here open up an alternative way, based on network theory, to analyze multilayer temporal configurations. Moreover, we show that the ABCD parameters can be used to compute the dispersion diagram ($\omega$ vs $k$) of time crystals.
@@ I Introduction
Maxwell’s equations describe with precision any electric and magnetic phenomena, or the interaction between both, involving classical field theory [1]. This includes the understanding and control of electromagnetic (EM) wave propagation, radiation, wave guiding, light-matter interactions at the human and nano scales, the production of basic electronic components such as resistors, inductors, capacitors or transformers, to name a few, the use of metamaterials, and many other interesting applications [2, 3, 4, 5, 6]. Certainly, Maxwell’s equations have transcended and transformed our society.
@@
In network analysis, a microwave or photonic network can be defined as a $N$-port system, whose inputs and outputs are interconnected via a governing matrix [10]. Impedance $[Z]$ (admittance $[Y]$) matrices are widely used solutions to relate the input voltages (currents) and output currents (voltages) in the network. Similarly, the scattering matrix $[S]$ does so with incident and reflected power waves. However, in many practical scenarios, the number of ports is reduced to $N=2$ (single input and single output). In this context, the use of ABCD parameters becomes particularly relevant.
@@
From a microwave theory perspective, the ABCD-matrix system can be seen as a “parameter black box” that links the input and output of a certain temporal system. In fact, treating a complex system as a black box is a very powerful tool that has been used historically in the context of engineering and, more specifically, in the electronics, microwave and antenna communities. This has been the case for the analysis and design of conventional time-invariant microwave components such as filters, amplifiers, power dividers, mixers, transmission lines, or general waveguiding devices [10]. Furthermore, the use of time modulation can be exploited in many engineering contexts, providing extra degrees of freedom and enhanced functionalities for applications such as MIMO communications or multi-target radar detection [18].
Despite the well-known foundations of ABCD parameters and their vast application in time-invariant engineering scenarios, the formalization of an analogous framework suited to time-varying EM problems is an open and evolving area of study [19, 20, 21]. In this work, we exploit microwave theory concepts and the temporal boundary conditions of the electric displacement field $\mathbf{D}$ and magnetic flux density field $\mathbf{B}$ to rearrange the input and output of the two-port time-varying system in a way that the resulting transfer matrix can be cascaded in a similar manner to conventional ABCD parameters. Then, we extract the ABCD parameters of a temporal transmission line (a temporal slab), and connect them with the associated scattering parameters and the computation of the dispersion diagram ($\omega$ vs $k$) in discrete and continuous time crystals. This formalism offers advantages compared to other matrix-based methodologies reported in the literature for time varying slabs [22, 23, 24], specially due to the substantial reduction of the number of matrices needed to compute the electromagnetic response of a time multi-layer system. Furthermore, the formalism is useful to provide valuable and specific physical insight to understand energetic and momentum processes in time-varying slabs.
@@ II Temporal Transfer ABCD Matrix
Figure 1: Use of transfer ABCD parameters in multilayer spatial and temporal interfaces. In both scenarios, ABCD matrices can be used to connect the EM fields at the input and output of the system. By exploiting the continuity of tangential $E_{t}$$H_{t}$} (total {$D$, BB}) fields in the spatial (temporal) problems, the fields can be rearranged such that the ABCD matrix of each layer can be multiplied to obtain the global response of the time-invariant (time-varying) system. The multilayer composite is formed by $N+2$ layers: the input layer [$(i)$, $n=0$], NN intermediate layers, and the output layer [$(o)$, $n=N+1$].
@@
As mentioned above, the ABCD-matrix formalism has proven to be quite efficient in characterizing cascade connections. As it is well known [10], the propagation of waves through a spatial dielectric layer of thickness $d_{n}$ admits to be represented by the following matrix:
@@
where $D_{n-1},B_{n-1},D_{n},B_{n}$ are the electric/magnetic flux vectors in the $n$th-order interface. In the following, we use the prime notation () to identify the ABCD parameters and avoid confusion with vectors $D$ and $B$. As we will see below, the result of the matrix allows us to indentify a temporal transmission line.
@@ II-A Temporal Transmission Line (Temporal Slab)
Let us consider the scenario depicted in the right panel of Fig. 1. In the scenario, a certain medium or material changes its electrical (or magnetic) properties over time at $t=t_{n-1}$. After a short period, the medium changes again its properties at $t=t_{n}$. The temporal region between $t_{n-1}$ and $t_{n}$ defines a temporal slab, a homogeneous medium that covers the whole space and is present for a limited time. Analogously to a spatial slab, a temporal slab is defined by its temporal length $T=t_{n}-t_{n-1}$, as well as its permittivity $\varepsilon_{n}$, permeability $\mu_{n}$ and the associated wave impedance $Z_{n}$. Please note that the wave impedance is defined as $Z_{n}=B_{n}/D_{n}$ in the $n$-th temporal slab. This definition results from a proper combination between the impedance definition $Z_{n}=E_{n}/H_{n}=\sqrt{\mu_{n}/\varepsilon_{n}}$, and the constitutive relations $D_{n}=\varepsilon_{n}E_{n}$, $B_{n}=\mu_{n}H_{n}$ .
A temporal slab can be formally described as a temporal transmission line, in analogy to the conventional spatial transmission line [10]. A temporal transmission line is an object that connects two time instants $t_{1}$ and $t_{2}$. For simplicity, we consider $t_{1}=0$ and $t_{2}=T$. Without loss of generality, the $D$ and $B$ fields can be represented at the input of the temporal transmission line in terms of forward (++) and backward (-) waves,
@@
Since we work with linear, isotropic and non-dispersive media, $D(t)=\varepsilon(t)E(t)$ and $B(t)=\mu(t)H(t)$ must hold. Thus, the coefficients $^{+}, DD^{-}$, $B^{+}$ and $B^{-}$ are related as follows
@@
After replacing eq. (6) into the expression for $D_{1}$ in (3), we realize that fields $D_{1}$, $D_{2}$ and $B_{2}$ are related as
@@
Similarly, the replacement of eq. (6) into the expression for $B_{1}$ leads to
@@
The connection between the fields at the two ports, input (1) and output (2), of the temporal transmission line can be better rearranged into matrix form:
@@
thus revealing that $A^{\prime}=\cos(\omega T)$, $B^{\prime}=-\mathrm{j}\frac{1}{Z}\sin(\omega T)$, $C^{\prime}=-\mathrm{j}Z\sin(\omega T)$, and $D^{\prime}=\cos(\omega T)$ are the transfer ABCD parameters of the temporal transmission line (temporal slab). Note that the frequency $\omega$ in eq. (9) varies across temporal slabs, while $k$ remains invariant. In general, the frequency in the $n$-th slab is given by $\omega=k/\sqrt{\varepsilon_{n}\mu_{n}}$.
It can be rapidly checked that the network is reciprocal ($A^{\prime}D^{\prime}-B^{\prime}C^{\prime}=1$) and symmetrical ($A^{\prime}=D^{\prime}$) in terms of electric and magnetic flux densities $D,B$. This temporal or momentum reciprocity does not necessarily conduct to the reciprocity for $E,H$ (spatial reciprocity), and vice versa. In addition, it is worth highlighting the units of the four ABCD parameters: $A^{\prime}$ [dimensionless], $B^{\prime}$ $\Omega^{-1}$], $C^{\prime}$ $\Omega$], and $D^{\prime}$ [dimensionless]. This represents a difference with other matrix formalisms that describe time-varying systems [22, 24].
@@
In the ABCD format, the number of matrices to multiply coincides with the number of temporal slabs $N$. Notice that this entails significant differences with other matrix formalisms reported in the literature, such as those in [22, 24]. There, the authors present a sound matrix approach that connects forward and backward $E$ waves, in a similar fashion to using scattering parameters, instead of directly working with $D$ and $B$. Since $E$ is not continuous across time interfaces, additional matrices (called matching/interface matrices) must be added for the entire calculation. For a temporal multilayer made of $N$ slabs, these methods typically need to employ $2N+1$ matrices, in contrast to the method here presented.
@@ II-B Associated Scattering Parameters
The analogy between spatial and temporal problems gives rise to the possibility of extracting the S-parameters associated with the temporal ABCD matrix. Now, the formal definition of the scattering matrix (S-matrix) takes the $D^{\pm}$, forward (++) and backward (-), waves at the input (1) and output (2) media. This contrasts to the spatial case, based on electric-field counterparts $E^{\pm}$. In both cases, however, there is a direct correspondence between them thanks to the constitutive relation $D^{\pm}=\varepsilon E^{\pm}$. The temporal S-matrix can be defined as:
$\begin{pmatrix}D_{2}^{-}\\ D_{2}^{+}\end{pmatrix}=\begin{pmatrix}S^{\prime}_{11}&S^{\prime}_{12}\\ S^{\prime}_{21}&S^{\prime}_{22}\end{pmatrix}\begin{pmatrix}D_{1}^{+}\\ D_{1}^{-}\end{pmatrix}\,$, (10)
@@
$S^{\prime}_{11}$ $=\triangle[-A^{\prime}-B^{\prime}Z_{0}+C^{\prime}/Z_{0}+D^{\prime}]$ (11)
$S^{\prime}_{12}$ $=\triangle[-A^{\prime}+B^{\prime}Z_{0}-C^{\prime}/Z_{0}+D^{\prime}]$ (12)
$S^{\prime}_{21}$ $=\triangle[A^{\prime}-B^{\prime}Z_{0}-C^{\prime}/Z_{0}+D^{\prime}]$ (13)
$S^{\prime}_{22}$ $=\triangle[A^{\prime}+B^{\prime}Z_{0}+C^{\prime}/Z_{0}+D^{\prime}]$ (14)
with $\triangle=[2A^{\prime}D^{\prime}-2B^{\prime}C^{\prime}]^{-1}$ and $Z_{0}=Z^{(i)}=Z^{(o)}$. We use the prime notation to identify the scattering parameters in the temporal problem and thus distinguish them from conventional time-invariant scenarios.
In the above definition, the parameters $S^{\prime}_{11}$ ($S^{\prime}_{21}$) would represent the ratio of the backward (forward) $D$-waves at the output port (2) and an incident forward $D$-wave at the input port (1). Similar physical reasoning is used to define $S^{\prime}_{12}$ and $S^{\prime}_{22}$, but with respect to an incident backward wave at the input port (1).
By assuming identical input/output media, the corresponding S-matrix for the electric field coincides with those in (11)-(14),
$S^{\prime}_{ij}=\frac{D_{2}^{\pm}}{D_{1}^{\pm}}=\frac{E_{2}^{\pm}}{E_{1}^{\pm}}\,.$ (15)
The definition of a more general framework of relationships involving different input/output media, $Z^{(i)}\neq Z^{(o)}$, is still under discussion and their physical implications are now focus of research.
@@ II-C Considerations about Conservation of Momentum Density
$E(z^{(i)})\cdot[H(z^{(i)})]^{*}=\underset{\text{Intermediate layers}}{\cdots}=E(z^{(o)})\cdot[H(z^{(o)})]^{*}\,$, (17)
@@
$D(t^{(i)})\cdot[B(t^{(i)})]^{*}=\underset{\text{Intermediate layers}}{\cdots}=D(t^{(o)})\cdot[B(t^{(o)})]^{*}\,.$ (19)
@@ II-D Time Crystal: Infinite Temporal Transmission Lines
with $a=0.5(\pm\sqrt{A^{\prime 2}-2A^{\prime}D^{\prime}+4B^{\prime}C^{\prime}+D^{\prime 2}}+A^{\prime}+D^{\prime})$. Eq. (22) offers a pathway to compute the dispersion diagram, $\omega$ vs $k$, in time crystals. In the above, note that $A^{\prime}D^{\prime}-B^{\prime}C^{\prime}=1$. The wavenumber dependence in eq. (22) is in the auxiliary parameter $a=a(k)$, since the ABCD parameters are dependent on the selected $k$ as well.
The advantage of the approach of the present paper is the possibility of extracting analytical expressions to compute $\omega(k)$ diagrams. Other methodologies in the literature, such as that reported in [26], address the problem from a numerical perspective by using complex root-finding algorithms. Our methodology provides formal analytical expressions for $\omega(k)$, including complex-valued frequency solutions. This is a significant advantage since evanescent- and, especially, growing-field solutions manifested via the imaginary part of the frequency can not be lightly disregarded in time-varying systems [27].
@@ III-A Temporal multilayer slab
Figure 2: Scattering parameters of an equally travel-distance multi-layered temporal structure with identical input and output media. $N=4$, $Z_{1}=Z_{0}/9$, $Z_{2}=Z_{0}/3$, and $Z^{(i)}=Z^{(o)}=Z_{0}$. Frequencies are normalized to the central frequency $f_{0}$.
@@ III-B Time Crystal
Figure 3: Dispersion diagram, computed with the ABCD formalism (white dotted lines) and the exact numerical solution (colored plot), for a time crystal given by the temporal modulation $\varepsilon_{r}(t)=\varepsilon_{r0}\left[1+m\cos(\Omega t)\right]$. (a) Discretization of the continuous temporal modulation (blue line) into NN steps (black lines). Dispersion diagram in the cases: (b) $N=5$, (c) $N=30$. Parameters: $\varepsilon_{r0}=1$, $T=0.1$ ns, $m=0.7$, and $\mu_{r}=1$.
@@
Let us consider a continous time crystal given by the time-periodic modulation $\varepsilon_{r}(t)=\varepsilon_{r0}\left[1+m\cos(\Omega t)\right]$, with $\varepsilon_{r0}$ being the DC component of the time modulation, $\Omega=2\pi/T$ ($T$ is the temporal period of the time crystal), and $m$ is the modulation factor. For simplicity, we only consider a modulation of $\varepsilon_{r}(t)$ ($\mu_{r}=1$), although a temporal magnetic response via $\mu_{r}(t)$ can be straightforwardly added. This type of time crystals has been successfully studied by Peter Halevi et al. [28, 26, 29]. By taking advantage of the periodic nature of $\varepsilon_{r}(t)$ and $\mu_{r}(t)$, the authors expand the electrical parameters into complex Fourier series and solve the problem in the frequency domain. Alternatively, we can solve the problem by discretizing $\varepsilon_{r}(t)$ into $N$ small pieces and the subsequent use of eq. (22) to compute the dispersion diagram. Thus, each piece represents a short temporal transmission line of length $T/N$ and impedance $Z_{n}=\sqrt{\mu_{r}(nT)/\varepsilon_{r}(nT)}$, which is modeled via the transfer ABCD matrix in eq. (9).
@@
Finally, it is important to remark that the present method allows to compute both the real and imaginary parts of the frequency. In time crystals, imaginary frequency components refer to time decay/grow of the corresponding fields. By following the engineering notation ($\mathrm{e}^{+j\omega t}\,\mathrm{e}^{-jkz}$), positive imaginary frequencies in the dispersion diagram imply attenuation over time, while negative imaginary frequencies imply amplification. These solutions are typical from the bandgap regions, manifesting from $k/K\in[0.4,0.55]$ in Fig. 3(b)-(c). Moreover, it is expected that the present ABCD approach can be extended to the analysis of uniform-velocity space-time crystals, in a similar fashion to was carried out in [30].